# Soot Planets instead of Water Worlds

### v1

@@ Abstract
Some low-density exoplanets are thought to be water-rich worlds that formed beyond the snow line of their protoplanetary disc, possibly accreting coequal portions of rock and water (R. Luque & E. Pallé, 2022; B. Bitsch et al., 2019b; A. Izidoro et al., 2021). However, the compositions of bodies within the Solar System and the stability of volatile-rich solids in accretionary disks suggest that a planet rich in water should also acquire as much as 40% refractory organic carbon (“soot”) (J. Li et al., 2021; E. A. Bergin et al., 2023). This would reduce the water mass fraction well below 50%, making the composition of these planets similar to those of Solar System comets (M. Rubin et al., 2019). Here we show that soot-rich planets, with or without water, can account for the low average densities of exoplanets that were previously attributed to a binary combination of rock and water. Formed in locations beyond the soot and/or snow lines in disks, these planets are likely common in our galaxy and already observed by JWST. The surfaces and interiors of soot-rich planets will be influenced by the chemical and physical properties of carbonaceous phases, and the atmospheres of such planets may contain plentiful methane and other hydrocarbons, with implications for photochemical haze generation and habitability.
@@ 1 Introduction
The masses and radii of the most common planets in the observable galaxy are intermediate between Earth and Neptune (W. J. Borucki et al., 2011; B. J. Fulton et al., 2017). These “sub-Neptune” exoplanets are typically less dense than the Earth (e.g., D. Charbonneau et al., 2009; R. Luque & E. Pallé, 2022), indicating that they contain abundant lighter components. Although such planets can be explained by rocky cores surrounded by large H2–He envelopes (J. E. Owen & Y. Wu, 2017; J. G. Rogers et al., 2023), they have alternatively been considered as “water worlds”, with as much as 50% H2O by mass (e.g., L. A. Rogers & S. Seager, 2010; R. Luque & E. Pallé, 2022). Formation of water worlds are thought to be a natural consequence of planetary accretion outside the “ice line” or “snow line”, where water ice is abundant (A. Morbidelli et al., 2000; B. Bitsch et al., 2019b), and therefore bodies with both rock and water ice should readily form. The possible existence of such sub-Neptune planets, which have been proposed to form a distinct population (R. Luque & E. Pallé, 2022), has elicited considerable inquiry, including the potential of temperate water worlds hosting liquid oceans (E. S. Kite & E. B. Ford, 2018; N. Madhusudhan et al., 2021).
@@
An alternative explanation for low-density sub-Neptunes, however, is that they are rich in carbonaceous materials, with or without appreciable water fractions. The conventional condensation sequence would not predict any appreciable carbon inside of the CO snow line, at about 30 AU (K. Lodders, 2003). The origin of the carbon phases found in primitive meteorites (and Ryugu, Bennu, etc.) is debated, but it is clear that this relatively refractory material was abundant in small primitive objects in the early solar system and was available to accrete into planetesimals and therefore planets. Carbon-rich grains may be inherited from the interstellar medium or forged in protoplanetary disk atmospheres. As posited in J. Li et al. (2021), the distribution and state of carbon in protoplanetary disks inside of the CO snow line is governed by irreversible sublimation instead of condensation. The “soot line” or “tar line” refers to the location where carbon-rich grains are destroyed via thermally driven reactions (M. E. Kress et al., 2010; K. Lodders, 2004; J. Li et al., 2021). It follows that planets accreted in the inner disk and beyond the soot line would acquire a large amount of carbon in the form of refractory organics (Fig. 1).
@@
Figure 1: Illustration of the soot line and water ice line structure of protoplanetary disks, and the three chemically distinct planet archetypes. The water ice line is also known as the water snow line or snow line. Compositions of the model rocky planet, soot planet, and soot-water worlds are shown as pie charts, with red, yellow, and blue segments representing the mass fractions of rock, soot, and water, respectively (Table A1). For the soot-water worlds, the pie chart and numbers refer to the end-member dry case and the numbers for the end-member wet case are: 36% rock, 14% soot, and 50% water. The soot line is defined by its sublimation temperature of 500 K under disk conditions (J. Li et al., 2021). The water ice line is placed at 150 K (M. Minissale et al., 2022). The locations of the soot lines and ice lines vary with time and across stellar mass range. For this study, it is the relative locations that matter. Figure credit Ari Gea and Sayo Studio.
@@
Indeed, protoplanetary disks are known to have a significant reservoir of refractory carbonaceous grains, or “soot”, which refers to a chemical component consisting of many phases, including meteoritic insoluble organic matter (IOM) and carbonaceous materials found in comets (M. E. Kress et al., 2010). It is a term that describes solids rich in carbon (C), hydrogen (H), oxygen (O), and with appreciable nitrogen (N), collectively known as CHON. The specific carriers of the carbon may include a variety of species of refractory organics such as refractory polyaromatic organics and semi-volatile organics found at the surface of comet 67P/CG (E. Quirico et al., 2016), refractory organics in the carbon-rich dust grains of comet 67P/CG and 1P/Halley (E. A. Bergin et al., 2015; M. Rubin et al., 2019, and references therein), and organic carbon carriers in ordinary chondrites (M. M. Grady & C. T. Pillinger, 1989) and martian meteorites (M. A. Sephton et al., 2003). Soot is stable and remains in the solid state to much greater temperatures ($\sim$500 K) (E. A. Bergin et al., 2015; H.-P. Gail & M. Trieloff, 2017; J. Li et al., 2021) than water ice ($\sim$150 K) (K. Lodders, 2003). A significant soot component would therefore be present in the formation zone of planets accreting outside of the snow line, unless there were significant periods of high temperature events driven by time variable accretion. The compositions of comets indicate that soot may make up to 40% of solid mass in this region (A. Bardyn et al., 2017; M. Rubin et al., 2019), while Uranus and Neptune are believed to be dominated by hydrocarbons rather than H2O (U. Malamud et al., 2024). Finally, the high C/O ratio of Jupiter observed by Galileo may reflect the accretion of abundant refractory carbon (K. Lodders, 2004).
@@
Importantly, because refractory carbon survives to higher temperatures than water ice, there are also likely regions where planets accrete both rock and soot, but little water (Fig. 1). Inside of the soot line, only rocky solids would be present. As a result, accretion of smaller gas-poor planets can potentially produce three generic archetypes depending on where accretion occurs: Rock-rich worlds with low carbon+water content (e.g., terrestrial planets), carbon-rich rocky worlds with low water content (E. A. Bergin et al., 2023), and rock/carbon/water worlds. In this study, we test the hypothesis of soot-rich planets against measured mass-radius relations. Furthermore, we discuss atmospheric characteristics for future detection of soot-rich planets and explore the implications of soot for planetary habitability.
@@ 2 Planet Models
We constructed models of exoplanet composition based on proto-solar elemental abundances and the distribution of solid materials found in comets (Fig. 1), as described in detail in Appendix. With the primary building blocks formed inside of the soot line, “rocky planets” are assumed to be composed entirely of metal and silicate, analogous to Earth. “Soot planets”, formed between the soot line and snow line, are made of both rock and soot. Planets forming outside of the water snow line are termed “soot-water worlds” and contain rock, soot, and water. We note that ‘soot-water worlds’ include a significant component of hydrocarbon-rich material and are therefore distinct from the ‘water worlds’ posited in earlier work (R. Luque & E. Pallé, 2022; B. Bitsch et al., 2019b; J. G. Rogers et al., 2023).
@@
Computing the exact mass-radius relationships of super-Earths or sub-Neptunes requires detailed and robust knowledge of the relevant phases at the high pressures-temperature conditions of planetary interiors, which is still largely lacking. In particular, predicting the fate of soot in planets has significant uncertainties because the prevailing carbonaceous phases for a given set of planetary parameters are poorly constrained. To overcome this challenge, we choose to treat the soot component as a fictive phase contributing to the properties of the planets, as described in Appendix. Using studies of meteorites (C. M. O. Alexander et al., 2012; E. A. Bergin et al., 2023), the ”soot” composition is set at C:H:O = 100:78:17 in atomic ratio. We estimated the density of soot at the reference condition of 1 bar and 298 K by examining the correlation between the density and mean atomic number for a diverse range of carbon-bearing phases, without considering its thermodynamic stability. The thermoelastic properties of soot are bounded by the highly compressible water ice and the highly incompressible diamond. These simplifying assumptions allow us to bound the M-R relations of soot planets, for comparisons with exoplanets with well-measured masses and radii (defined here as uncertainties less than 20% and 10%, respectively).
@@
We consider two end-member model planets that are either fully stratified or fully mixed. A multi-layer stratified planet consists of a metallic core enclosed in a silicate mantle, and where present, overlain by a hydrocarbon-rich layer, and finally a water ice surface layer. A single-layer mixed planet represents an extreme case where iron, silicate, soot, and ice are fully mixed at all depths. It is usually assumed that planets have sufficient energy to differentiate and form layered interiors. However, the high temperatures prevailing in sub-Neptunes may produce exotic chemistry in which metal, silicate and water are fully miscible(E. D. Young et al., 2025, 2024). We expect real planets to lie in between these two extremes, because water, soot, silicate, and metal may be partially or wholly miscible. Details of the mass-radius (M-R) calculation are described in the Appendix.
@@ 3 Results and Discussion
Our composition models (Table A1) integrated astronomical and cosmochemical constraints on solar composition (K. Lodders, 2003), the disposition of carbon during planet formation (J. Li et al., 2021), compositions of insoluble organic matter (IOM) in chondrites (C. Alexander et al., 2017), and the dust-to-ice ratios in cometary materials (M. Rubin et al., 2019).
@@
The mass-radius relations of the model planets are established and compared to the collection of known exoplanets (Fig. 3). For the model soot planet and soot-water worlds, a range of radii are possible at a given mass, reflecting the limits we assume on whether the soot component is highly incompressible as diamond or highly compressible as H2O ice (Table A3, Fig. A3). At a given composition and mass, the calculated radius of a differentiated planet is found to be smaller than its undifferentiated counterpart. The difference would be at least partially offset by thermal expansion associated with differentiation. Exoplanets of $\lesssim$10 Earth masses exhibit a wide range of densities, with many observed mini-Neptunes following similar mass-radius relationships as soot planets or soot-water worlds. The predicted M-R relations for the realistic water worlds, which incorporates soot, is similar to that predicted previously for a 50% water planet with no carbon R. Luque & E. Pallé (2022). Thus, based on the expected M-R relations alone, observed exoplanets previously interpreted as rock+water planets cannot be distinguished from water-rich planets incorporating significant soot. This suggests that a major compositional component of such planets may be overlooked.
@@
Previous studies also considered carbon-bearing planets, but they arise from different implied formation scenarios that do not account for refractory carbon in the protoplanetary disk, and therefore they differ from our soot planets in fundamental aspects. For example, the vast majority of stellar C/O ratios for known planetary systems mostly fall between 0.3 and 0.8, with the medium value at the solar value of 0.5 (Fig. A1). Carbon-rich exoplanets have been proposed for stars with C/O ratios that are close to or exceed unity. (e.g.,(N. Madhusudhan et al., 2012; F. Miozzi et al., 2018; K. Hakim et al., 2019)), In contrast, our model soot planets have sub-solar C/O ratios ranging from one-half to one-sixth of the solar value. Moreover, other studies that considered carbon-rich exoplanet atmospheres have typically invoked ice-line arguments as the source of the augmented carbon, whereas our soot-rich planets formed over a wide range of the accretion disc beyond the soot line and contain a substantial hydrocarbon-rich component (e.g., N. Madhusudhan et al., 2011; K. I. Öberg et al., 2011a; M. Ali-Dib et al., 2014). Recently, B. Peng & D. Valencia (2024) explored the potential for carbon-rich mantles leading to the creation of Venus analogs with thick, carbon-dominated atmospheres. They adopted the CI level of carbon content for planets forming inside the water ice line and did not discuss the source of carbon in the formation process.
@@
The M-R relations shown in Fig. 3 have substantial degeneracy (e.g., S. Seager et al., 2007; L. A. Rogers & S. Seager, 2010). In this work, we identify new degeneracies that should be considered and that challenge the water world interpretation. Soot-rich planets or cores of sub-Neptunes (with or without substantive water inventories) are predicted by planet formation models that explain key astronomical and cosmochemical observations. They also adequately explain the observed mass-radius relations of low-mass low-density exoplanets. Therefore, we must contend with this level of compositional complexity, especially since the mass fractions of soot and water may vary depending on the their formation path (F. Binkert & T. Birnstiel, 2023). The model planets illustrated here are based on the range of observed dust to ice ratios in cometary material (M. Rubin et al., 2019). This ratio may vary in part because a planet may lose volatiles during accretion via parent-body processing (T. Lichtenberg et al., 2023) and impact degassing (H. E. Schlichting & S. Mukhopadhyay, 2018).
@@
Our model planets do not contain gaseous (H2+He) envelopes. With increasing mass, planets are more likely to attract and retain appreciable nebular hydrogen (J. E. Owen & Y. Wu, 2013). Therefore, H2-rich atmospheres are quite likely present in massive planets, but are less likely to be preserved in low mass planets. For more massive planets, a gaseous envelope may contribute to the mass - radius relations of observed low mean density planets. For example, a nebular envelope of 1% planet mass increases the radius of a rocky planet by 20 to 30% (Fig. 3). Therefore, relatively massive ($\gtrsim 1$ $M_{\oplus}$; (J. Owen et al., 2020)) soot planets or soot-water worlds cannot be distinguished from rocky planets with H2 envelopes based on M-R relationships alone. Nevertheless, we expect larger planets with hydrogen envelopes to have soot-rich mantles that will likely alter the atmospheric composition (E. A. Bergin et al., 2023). Smaller planets and those close to their star are less likely to have retained their primary gaseous envelopes (J. Owen et al., 2020), so soot or soot-water worlds are more likely explanations for these low mass, low density planets. A key point is that no single compositional scenario describes the full set of observed exoplanets, so considerable compositional diversity (i.e., different admixtures of iron, rock, water, hydrogen, and soot) are needed to explain the full range of planetary radii observed at a given mass, with massive hydrogen envelopes being more likely at higher masses and above the so-called ‘radius valley’ (B. J. Fulton et al., 2017).
@@
The distribution of carbonaceous phases in a soot-rich planet has been little-explored. Major possible sinks for carbon in known planets from the Solar System include the metallic cores of differentiated terrestrial planets R. A. Fischer et al. (2020); J. Li et al. (2021) and clathrate ice layers in the ice giants Uranus and Neptune (U. Malamud et al., 2024). Metallic cores of planets can only store a small fraction of carbon delivered as soot: The maximum concentration of carbon in an iron-rich core is below 10 wt% R. Dasgupta & D. Walker (2008), corresponding to 2.5 wt.% C in a planet with 25 wt.% metallic iron, which is much less than the average C concentration of a soot planet (Table A1). Clathrate ices are plausible for planets with cold interiors J. I. Lunine & D. J. Stevenson (1985), but not for those more thermally vigorous.
@@
Soot-rich planets are expected to possess distinct atmospheric compositions. Thermal processing of soot-rich components in planetary interiors, either in isolated layers or mixed with silicate, is expected to yield refractory phases such as graphite, diamond, or carbide (V. A. Davydov et al., 2004). =Through interactions with H2-rich envelopes, the mantles would be reduced with high H/C ratios. Because the solubility of C-H-rich components in silicate liquid and minerals is limited (L. S. Armstrong et al., 2015; P. Ardia et al., 2013), significant fractions of methane and other simple hydrocarbons are expected to be released from the interior. A resulting methane-rich atmosphere may naturally lead to the formation of hydrocarbon hazes, akin to the tholins in Titan’s atmosphere (E. Miller-Ricci Kempton et al., 2012; C. V. Morley et al., 2013; E. A. Bergin et al., 2023). Indeed, many sub-Neptunes have produced featureless spectra indicating the likely presence of clouds or photochemical haze, perhaps supporting our concept of a soot planet (e.g., L. Kreidberg et al., 2014; J. Brande et al., 2024).
@@
JWST observations of sub-Neptunes have recently allowed for the direct detection of carbon-bearing species in their atmospheres (N. Madhusudhan et al., 2023; M. Holmberg & N. Madhusudhan, 2024; B. Benneke et al., 2024). An active area of research attempts to tie the observed atmospheric composition to the properties of a planet’s bulk interior (e.g., M. C. Nixon et al., 2024; N. F. Wogan et al., 2024; F. E. Rigby et al., 2024). Presently it is unclear whether the observed composition of any individual sub-Neptune atmosphere indicates a significant fraction of carbon in its bulk, but tentative evidence supports our definition of either a soot planet or a soot-water world. For example, the discovery of abundant CO2 and CH4 in the atmospheres of K2-18 b and TOI-270 d, accompanied by H2O in the latter case (N. Madhusudhan et al., 2023; M. Holmberg & N. Madhusudhan, 2024; B. Benneke et al., 2024), indicates a reservoir of carbon participating in atmospheric chemistry. Linking the atmospheric carbon-to-oxygen ratio (C/O) to the carbon content of the planet’s bulk interior requires models of the interface between the deep atmosphere and an underlying magma ocean, which will be the subject of future work. In the absence of a considerable soot component, one would qualitatively expect a very low atmospheric C/O because water outgasses readily from silicates in contact with hydrogen-dominated atmospheres, but we are not aware of a comparable mechanism to release abundant carbon-rich gases from an initially carbon-poor interior (I. Blanchard et al., 2022; C. Seo et al., 2024). Thus the moderately high C/O for the sub-Neptune TOI-270 d ($\sim$0.5) may indicate that it is a soot planet.

### v2

@@ Abstract
Some low-density exoplanets are thought to be water-rich worlds that formed beyond the snow line of their protoplanetary disc, possibly accreting coequal portions of rock and water (Luque22; Bitsch2019_waterworld; Izidoro2021). However, the compositions of bodies within the Solar System and the stability of volatile-rich solids in accretionary disks suggest that a planet rich in water should also acquire as much as 40% refractory organic carbon (“soot”) (Li21; Bergin2023). This would reduce the water mass fraction well below 50%, making the composition of these planets similar to those of Solar System comets (Rubin19). Here we show that soot-rich planets, with or without water, can account for the low average densities of exoplanets that were previously attributed to a binary combination of rock and water. Formed in locations beyond the soot and/or snow lines in disks, these planets are likely common in our galaxy and already observed by JWST. The surfaces and interiors of soot-rich planets will be influenced by the chemical and physical properties of carbonaceous phases, and the atmospheres of such planets may contain plentiful methane and other hydrocarbons, with implications for photochemical haze generation and habitability.
@@ 1 Introduction
The masses and radii of the most common planets in the observable galaxy are intermediate between Earth and Neptune (Borucki2011; fulton17). These “sub-Neptune” exoplanets are typically less dense than the Earth (e.g., Charbonneau2009; Luque22), indicating that they contain abundant lighter components. Although such planets can be explained by rocky cores surrounded by large H2–He envelopes (owen17; rogers2023conclusive), they have alternatively been considered as “water worlds”, with as much as 50% H2O by mass (e.g., rogers10; Luque22). Formation of water worlds are thought to be a natural consequence of planetary accretion outside the “ice line” or “snow line”, where water ice is abundant (Mobidelli2000; Bitsch2019_waterworld), and therefore bodies with both rock and water ice should readily form. The possible existence of such sub-Neptune planets, which have been proposed to form a distinct population (Luque22), has elicited considerable inquiry, including the potential of temperate water worlds hosting liquid oceans (kite2018habitability; madhu21).
@@
An alternative explanation for low-density sub-Neptunes, however, is that they are rich in carbonaceous materials, with or without appreciable water fractions. The conventional condensation sequence would not predict any appreciable carbon inside of the CO snow line, at about 30 AU (Lodders03). The origin of the carbon phases found in primitive meteorites (and Ryugu, Bennu, etc.) is debated, but it is clear that this relatively refractory material was abundant in small primitive objects in the early solar system and was available to accrete into planetesimals and therefore planets. Carbon-rich grains may be inherited from the interstellar medium or forged in protoplanetary disk atmospheres. As posited in Li21, the distribution and state of carbon in protoplanetary disks inside of the CO snow line is governed by irreversible sublimation instead of condensation. The “soot line” or “tar line” refers to the location where carbon-rich grains are destroyed via thermally driven reactions (Kress2010; Lodders2004; Li21). It follows that planets accreted in the inner disk and beyond the soot line would acquire a large amount of carbon in the form of refractory organics (Fig. 1).
@@
Figure 1: Illustration of the soot line and water ice line structure of protoplanetary disks, and the three chemically distinct planet archetypes. The water ice line is also known as the water snow line or snow line. Compositions of the model rocky planet, soot planet, and soot-water worlds are shown as pie charts, with red, yellow, and blue segments representing the mass fractions of rock, soot, and water, respectively (Table A1). For the soot-water worlds, the pie chart and numbers refer to the end-member dry case and the numbers for the end-member wet case are: 36% rock, 14% soot, and 50% water. The soot line is defined by its sublimation temperature of 500 K under disk conditions (Li21). The water ice line is placed at 160 K (Minissale22). The locations of the soot lines and ice lines vary with time and across stellar mass range. The soot line and ice line are drawn for illustrative purposes, Their actual shapes depend on the pressure distribution in the disk, as well as the intensity of viscous heating and the contribution from irradiation from the central star, which vary over time and with physical conditions in the disk. For this study, it is the relative locations that matter. Figure credit Ari Gea and Sayo Studio.
@@
Indeed, protoplanetary disks are known to have a significant reservoir of refractory carbonaceous grains, or “soot”, which refers to a chemical component consisting of many phases, including meteoritic insoluble organic matter (IOM) and carbonaceous materials found in comets (Kress2010). It is a term that describes solids rich in carbon (C), hydrogen (H), oxygen (O), and with appreciable nitrogen (N), collectively known as CHON. The specific carriers of the carbon may include a variety of species of refractory organics such as refractory polyaromatic organics and semi-volatile organics found at the surface of comet 67P/CG (2016Icar..272...32Q), refractory organics in the carbon-rich dust grains of comet 67P/CG and 1P/Halley (Bergin15; Rubin19, and references therein), and organic carbon carriers in ordinary chondrites (Grady89a) and martian meteorites (Sephton03). Soot is stable and remains in the solid state to much greater temperatures ($\sim$500 K) (Bergin15; Gail17; Li21) than water ice ($\sim$150 K) (Lodders03). A significant soot component would therefore be present in the formation zone of planets accreting outside of the snow line, unless there were significant periods of high temperature events driven by time variable accretion. The compositions of comets indicate that soot may make up to 40% of solid mass in this region (Bardyn2017; Rubin19), while Uranus and Neptune are believed to be dominated by hydrocarbons rather than H2O (Malamud2024). Finally, the high C/O ratio of Jupiter observed by Galileo may reflect the accretion of abundant refractory carbon (Lodders2004).
@@
Importantly, because refractory carbon survives to higher temperatures than water ice, there are also likely regions where planets accrete both rock and soot, but little water (Fig. 1). Inside of the soot line, only rocky solids would be present. As a result, accretion of smaller gas-poor planets can potentially produce three generic archetypes depending on where accretion occurs: Rock-rich worlds with low carbon+water content (e.g., terrestrial planets), carbon-rich rocky worlds with low water content (Bergin2023), and rock/carbon/water worlds. In this study, we test the hypothesis of soot-rich planets against measured mass-radius relations. Furthermore, we discuss atmospheric characteristics for future detection of soot-rich planets and explore the implications of soot for planetary habitability.
@@ 2 Planet Models
We constructed models of exoplanet composition based on proto-solar elemental abundances and the distribution of solid materials found in comets (Fig. 1), as described in detail in Appendix. With the primary building blocks formed inside of the soot line, “rocky planets” are assumed to be composed entirely of metal and silicate, analogous to Earth. “Soot planets”, formed between the soot line and snow line, are made of both rock and soot. Planets forming outside of the water snow line are termed “soot-water worlds” and contain rock, soot, and water. We note that ‘soot-water worlds’ include a significant component of hydrocarbon-rich material and are therefore distinct from the ‘water worlds’ posited in earlier work (Luque22; Bitsch2019_waterworld; rogers2023conclusive).
@@
Computing the exact mass-radius relationships of super-Earths or sub-Neptunes requires detailed and robust knowledge of the relevant phases at the high pressures-temperature conditions of planetary interiors, which is still largely lacking. In particular, predicting the fate of soot in planets has significant uncertainties because the prevailing carbonaceous phases for a given set of planetary parameters are poorly constrained. To overcome this challenge, we choose to treat the soot component as a fictive phase contributing to the properties of the planets, as described in Appendix. Using studies of meteorites (Alexander12; Bergin2023), the ”soot” composition is set at C:H:O = 100:78:17 in atomic ratio. We estimated the density of soot at the reference condition of 1 bar and 298 K by examining the correlation between the density and mean atomic number for a diverse range of carbon-bearing phases, without considering its thermodynamic stability. The thermoelastic properties of soot are bounded by the highly compressible water ice and the highly incompressible diamond. These simplifying assumptions allow us to bound the M-R relations of soot planets, for comparisons with exoplanets with well-measured masses and radii (defined here as uncertainties less than 20% and 10%, respectively).
@@
We consider two end-member model planets that are either fully stratified or fully mixed. A multi-layer stratified planet consists of a metallic core enclosed in a silicate mantle, and where present, overlain by a hydrocarbon-rich layer, and finally a water ice surface layer. A single-layer mixed planet represents an extreme case where iron, silicate, soot, and ice are fully mixed at all depths. It is usually assumed that planets have sufficient energy to differentiate and form layered interiors. However, the high temperatures prevailing in sub-Neptunes may produce exotic chemistry in which metal, silicate and water are fully miscible(Young2025arXiv250700947Y; young2024phase). We expect real planets to lie in between these two extremes, because water, soot, silicate, and metal may be partially or wholly miscible. Details of the mass-radius (M-R) calculation are described in the Appendix.
@@ 3 Results and Discussion
Our composition models (Table A1) integrated astronomical and cosmochemical constraints on solar composition (Lodders03), the disposition of carbon during planet formation (Li21), compositions of insoluble organic matter (IOM) in chondrites (Alexander17), and the dust-to-ice ratios in cometary materials (Rubin19).
@@
The mass-radius relations of the model planets are established and compared to the collection of known exoplanets (Fig. 3). For the model soot planet and soot-water worlds, a range of radii are possible at a given mass, reflecting the limits we assume on whether the soot component is highly incompressible as diamond or highly compressible as H2O ice (Table A3, Fig. A3). At a given composition and mass, the calculated radius of a differentiated planet is found to be smaller than its undifferentiated counterpart. The difference would be at least partially offset by thermal expansion associated with differentiation. Exoplanets of $\lesssim$10 Earth masses exhibit a wide range of densities, with many observed mini-Neptunes following similar mass-radius relationships as soot planets or soot-water worlds. The predicted M-R relations for the realistic water worlds, which incorporates soot, is similar to that predicted previously for a 50% water planet with no carbon Luque22. Thus, based on the expected M-R relations alone, observed exoplanets previously interpreted as rock+water planets cannot be distinguished from water-rich planets incorporating significant soot. This suggests that a major compositional component of such planets may be overlooked.
@@
Previous studies also considered carbon-bearing planets, but they arise from different implied formation scenarios that do not account for refractory carbon in the protoplanetary disk, and therefore they differ from our soot planets in fundamental aspects. For example, the vast majority of stellar C/O ratios for known planetary systems mostly fall between 0.3 and 0.8, with the medium value at the solar value of 0.5 (Fig. A1). Carbon-rich exoplanets have been proposed for stars with C/O ratios that are close to or exceed unity. (e.g.,(Madhu12; 2018JGRE..123.2295M; Hakim19)), In contrast, our model soot planets have sub-solar C/O ratios ranging from one-half to one-sixth of the solar value. Moreover, other studies that considered carbon-rich exoplanet atmospheres have typically invoked ice-line arguments as the source of the augmented carbon, whereas our soot-rich planets formed over a wide range of the accretion disc beyond the soot line and contain a substantial hydrocarbon-rich component (e.g., madhus11; oberg11_c2d; alidib14). Recently, Peng2024ApJ...976..202P explored the potential for carbon-rich mantles leading to the creation of Venus analogs with thick, carbon-dominated atmospheres. They adopted the CI level of carbon content for planets forming inside the water ice line and did not discuss the source of carbon in the formation process.
@@
The M-R relations shown in Fig. 3 have substantial degeneracy (e.g., Seager2007; rogers10). In this work, we identify new degeneracies that should be considered and that challenge the water world interpretation. Soot-rich planets or cores of sub-Neptunes (with or without substantive water inventories) are predicted by planet formation models that explain key astronomical and cosmochemical observations. They also adequately explain the observed mass-radius relations of low-mass low-density exoplanets. Therefore, we must contend with this level of compositional complexity, especially since the mass fractions of soot and water may vary depending on the their formation path (Binkert2023). The model planets illustrated here are based on the range of observed dust to ice ratios in cometary material (Rubin19). This ratio may vary in part because a planet may lose volatiles during accretion via parent-body processing (Lichtenberg2023) and impact degassing (2018SSRv..214...34S).
@@
Our model planets do not contain gaseous (H2+He) envelopes. With increasing mass, planets are more likely to attract and retain appreciable nebular hydrogen (Owen2013). Therefore, H2-rich atmospheres are quite likely present in massive planets, but are less likely to be preserved in low mass planets. For more massive planets, a gaseous envelope may contribute to the mass - radius relations of observed low mean density planets. For example, a nebular envelope of 1% planet mass increases the radius of a rocky planet by 20 to 30% (Fig. 3). Therefore, relatively massive ($\gtrsim 1$ $M_{\oplus}$; (owen2020hydrogen)) soot planets or soot-water worlds cannot be distinguished from rocky planets with H2 envelopes based on M-R relationships alone. Nevertheless, we expect larger planets with hydrogen envelopes to have soot-rich mantles that will likely alter the atmospheric composition (Bergin2023). Smaller planets and those close to their star are less likely to have retained their primary gaseous envelopes (owen2020hydrogen), so soot or soot-water worlds are more likely explanations for these low mass, low density planets. A key point is that no single compositional scenario describes the full set of observed exoplanets, so considerable compositional diversity (i.e., different admixtures of iron, rock, water, hydrogen, and soot) are needed to explain the full range of planetary radii observed at a given mass, with massive hydrogen envelopes being more likely at higher masses and above the so-called ‘radius valley’ (fulton17).
@@
The distribution of carbonaceous phases in a soot-rich planet has been little-explored. Major possible sinks for carbon in known planets from the Solar System include the metallic cores of differentiated terrestrial planets Fischer20; Li21 and clathrate ice layers in the ice giants Uranus and Neptune (Malamud2024). Metallic cores of planets can only store a small fraction of carbon delivered as soot: The maximum concentration of carbon in an iron-rich core is below 10 wt% Dasgupta_Walker_2008, corresponding to 2.5 wt.% C in a planet with 25 wt.% metallic iron, which is much less than the average C concentration of a soot planet (Table A1). Clathrate ices are plausible for planets with cold interiors Lunine_Stevenson_1985, but not for those more thermally vigorous.
@@
Soot-rich planets are expected to possess distinct atmospheric compositions. Thermal processing of soot-rich components in planetary interiors, either in isolated layers or mixed with silicate, is expected to yield refractory phases such as graphite, diamond, or carbide (Davydov2004). =Through interactions with H2-rich envelopes, the mantles would be reduced with high H/C ratios. Because the solubility of C-H-rich components in silicate liquid and minerals is limited (Armstrong15; Ardia13), significant fractions of methane and other simple hydrocarbons are expected to be released from the interior. A resulting methane-rich atmosphere may naturally lead to the formation of hydrocarbon hazes, akin to the tholins in Titan’s atmosphere (millerricci12; morley13; Bergin2023). Indeed, many sub-Neptunes have produced featureless spectra indicating the likely presence of clouds or photochemical haze, perhaps supporting our concept of a soot planet (e.g., Kreidberg14; brande24).
@@
JWST observations of sub-Neptunes have recently allowed for the direct detection of carbon-bearing species in their atmospheres (madhu23; holmberg24; benneke24). An active area of research attempts to tie the observed atmospheric composition to the properties of a planet’s bulk interior (e.g., nixon24; wogan24; rigby24). Presently it is unclear whether the observed composition of any individual sub-Neptune atmosphere indicates a significant fraction of carbon in its bulk, but tentative evidence supports our definition of either a soot planet or a soot-water world. For example, the discovery of abundant CO2 and CH4 in the atmospheres of K2-18 b and TOI-270 d, accompanied by H2O in the latter case (madhu23; holmberg24; benneke24), indicates a reservoir of carbon participating in atmospheric chemistry. Linking the atmospheric carbon-to-oxygen ratio (C/O) to the carbon content of the planet’s bulk interior requires models of the interface between the deep atmosphere and an underlying magma ocean, which will be the subject of future work. In the absence of a considerable soot component, one would qualitatively expect a very low atmospheric C/O because water outgasses readily from silicates in contact with hydrogen-dominated atmospheres, but we are not aware of a comparable mechanism to release abundant carbon-rich gases from an initially carbon-poor interior (blanchard22; seo24). Thus the moderately high C/O for the sub-Neptune TOI-270 d ($\sim$0.5) may indicate that it is a soot planet.